Two forces F1 and F2 are acting on an object F1 =80N acting at bearing of 45° and F2 =60N acting S 45° W

1
Expert's answer
2022-02-06T14:33:31-0500

"F=F_1+F_2=\frac{80\sqrt 2}2-\frac{60\sqrt2}2=14.3~N."